you have 28 coins, they are nickles and dimes. The value of the coins is $2.05
Define the variables, Write a system of equations for the solutions

Answer by checkley71(8403)

10X+5(28-X)=205 OR 10X+140-5X=205 OR 10X-5X=205-140 OR 5X=65 OR X=13
10X13=130 & 5(28-13) OR 5X15=65 SO 130+65=205 OR 13 DIMES PLUS 15 NICKLES=$2.05